# Align-GRAG: Reasoning-Guided Dual Alignment for Graph Retrieval-Augmented Generation

### v1

@@ Abstract
Large language models (LLMs) have demonstrated remarkable capabilities, but still struggle with issues like hallucinations and outdated information. Retrieval-augmented generation (RAG) addresses these issues by grounding LLM outputs in external knowledge with an Information Retrieval (IR) system. Building on this foundation, graph-based RAG systems go a step further by retrieving subgraphs, which preserve the relationships between knowledge entities and provide more comprehensive context. However, graph RAG faces two challenges: (1) Retrieving relevant information introduces irrelevant nodes (especially in dense graph databases, where retrieval usually extends to adjacent nodes), and leads to overly lengthy inputs that hinder efficiency; (2) The representation gap between graph and language during generation with LLMs limits the ability to fully leverage graph structures for enhanced understanding. To address these limitations, we propose Align-GRAG, a novel reasoning-guided dual alignment framework in post-retrieval phrase. It first formulates a subgraph by retrieving nodes and edges. Then an Aligner is proposed to jointly optimizes a graph encoder with LLM-summarized reasoning. It achieves dual alignment of graph node and representation by leveraging KL divergence loss and contrastive loss, facilitating efficient pruning of irrelevant knowledge and establishing a unified semantic space. The Generator integrates the aligned graph data with LLM to produce coherent and accurate answers. Experiments on GraphQA benchmark across three tasks (including common sense reasoning, scene graph understanding, and knowledge graph reasoning) validate the effectiveness of our method. The code will be available upon accepted.
1 Introduction
@@
Recent advancements in natural language processing have demonstrated the remarkable capabilities of large language models (LLMs), such as GPT [1] and Llama [2], in understanding, reasoning, and handling complex tasks [3, 4]. The massive parameterized knowledge encoded within LLMs has enabled them to excel in information retrieval [5, 6, 7]. Despite these advancements, concerns about the interpretability of LLM remain, leading to challenges like hallucinations (fabrication of false or misleading information) [8, 9] and reliance on outdated data [10]. These issues are especially concerning in high-stakes fields such as healthcare [11] and law [12].
@@
Retrieval-augmented generation (RAG) systems have been developed to solve these problems [10]. By integrating external information retrieval (IR) systems, RAG retrieves knowledge from external databases, ensuring that LLMs are equipped with up-to-date and relevant information tailored to the user’s query. By grounding the LLM’s outputs in verifiable, external knowledge, RAG systems significantly improve the reliability and accuracy of content generated by LLMs [13]. Nevertheless, in real-world scenarios, RAG systems often divide long context into independent chunks, overlooking the deeper connections between fragments and lacking a global perspective [14, 15]. Moreover, a significant portion of data inherently exhibits a graph-like structure, such as in recommendation systems [16], the Web, and knowledge graphs [17]. To address these, researchers have proposed graph-based RAG (GRAG) [14, 18, 19, 20] as an enhancement to traditional RAG methods. Instead of retrieving isolated text chunks where information is treated independently of its structural relationships, GRAG retrieves subgraphs from databases [18]. These subgraphs maintain the interconnected relationships between knowledge entities, providing richer context and enabling more accurate, contextually relevant responses.
@@
However, despite its potential, integrating RAG with graphs is non-trivial, presenting two key challenges: (1) Irrelevant Knowledge: The retrieval stage often introduces irrelevant information, hindering the LLM’s attention on useful knowledge (especially in dense graph databases, where retrieval usually extends to adjacent nodes), and the generated inputs become overly lengthy, further causing computational inefficiency. Some traditional RAG approaches [21, 22, 23] aim to tackle this challenge by incorporating a reranker in the post-retrieval stage [13, 10]. For instance, BGE [21] leverages cross-encoder-based architectures to effectively capture interactions between queries and documents, enabling more precise relevance estimation. However, in the GRAG setting, where chunks are typically represented as nodes or triplets, this approach can disrupt structural information and overlook the connections between knowledge. Meanwhile, existing GRAG methods [15] focus primarily on retrieval-stage optimizations (e.g., GraphRAG [14] builds community graphs for long documents to facilitate retrieval and summarization), resulting in relatively shallow similarity that lack deeper semantic matching capabilities. (2) Representation Gap: In the generation stage, when integrating LLMs with graph structure embeddings, such as those generated by Graph Neural Network (GNN), challenge arises due to the representation gap: graph embeddings are designed to capture structural properties, which misaligns with the sequence-based representations used by LLMs. This misalignment limits the ability to fully leverage graph for enhanced reasoning [24, 25]. In the field of LLMs, most existing GRAG methods [20, 26, 27, 28, 29, 30] primarily treat graph data as plain text inputs, lacking a dedicated encoder for structural information, which limits their understanding capabilities. Some methods [19, 31, 18] incorporate graph encoders, but typically rely on straightforward approaches like concatenating graph embeddings with LLM inputs. For instance, G-Retriever [18] uses a projector to map graph embeddings, while GNP [19] applies cross-modality pooling to fuse text and graph representations. However, neither approach explicitly aligns these distinct representation spaces, leaving the underlying gap unresolved. These research gaps present a question: Can we better prune irrelevant nodes while keeping structural information, and effectively encode graph representations for utilization by LLMs?
@@
In this work, we propose Align-GRAG, a novel reasoning-guided dual alignment framework specifically designed for GRAG. During the retrieval stage, we calculate the similarity between nodes and edges to extract the initial subgraph. To address the two key challenges in GRAG, we introduce an innovative graph aligner module for the post-retrieval pharse. We leverage LLM-based summarization ability with well-crafted prompts to generate reasoning chains that bridge information gaps. The aligner achieves dual alignment with reasoning chains as optimization target: (1) Graph Node Alignment, which ranks critical reasoning nodes, pruning irrelevant knowledge; and (2) Graph Representation Alignment, ensuring unified semantic space between graph and language representations via contrastive learning. This dual alignment refines subgraphs to focus on relevant knowledge, enabling the generator to produce accurate, context-aware responses grounded in the graph’s structure. Extensive experiments are conducted on GraphQA benchmark [18], covering tasks such as commonsense reasoning, scene graph understanding, and knowledge graph reasoning, which demonstrate the superior performance of Align-GRAG compared to other baselines and highlight the effectiveness of Aligner module. Overall, this work makes three key contributions:
@@
2 Preliminaries
@@
This section introduces key concepts and notations, including textual graphs and task formulation.
@@
Textual Graph. A textual graph is a graph where nodes and edges are enriched with textual information, capturing both structural and semantic details. Formally, a textual graph is defined as $\mathcal{G}=(\mathcal{V},\mathcal{E},\{t_{n}\}_{n\in\mathcal{V}},\{t_{e}\}_{e\in\mathcal{E}})$, where $\mathcal{V}$ and $\mathcal{E}$ are the sets of nodes and edges. $t_{n}\in\mathcal{D}^{L_{n}}$ represents the text associated with a node $n\in\mathcal{V}$, where $\mathcal{D}$ is the vocabulary, and $L_{n}$ is the text length. Similarly, $t_{e}\in\mathcal{D}^{L_{e}}$ is the text for an edge $e\in\mathcal{E}$, with $L_{e}$ as its length.
@@
Task Formulation. This work addresses RAG with textual graph. The goal is to retrieve relevant information from textual graphs and generate accurate responses. Given a query $t_{q}$, a sequence of tokens from $\mathcal{D}$, the model retrieves a subgraph $\mathcal{G}_{r}=(\mathcal{V}_{r},\mathcal{E}_{r})$ from $\mathcal{G}$ based on the semantic similarity of $t_{n}$, $t_{e}$ with $t_{q}$, where $\mathcal{V}_{r}\subseteq\mathcal{V}$ and $\mathcal{E}_{r}\subseteq\mathcal{E}$. The retrieved nodes, edges, and texts are refined to improve the input quality for the LLM. During generation, the retrieved subgraph, query $t_{q}$, and prompt $\mathcal{P}$ are provided as input to the LLM, which produces the final output $Y$, grounded in the retrieved knowledge and graph context.
3 Methodology
@@
In this section, we introduce Align-GRAG framework, which is composed of three components: Graph Retriever (Sec. 3.1), Graph Aligner (Sec. 3.2), and Graph Generator (Sec. 3.3) modules. The overall framework pipeline is depicted in Figure 1. In Figure 1, ➀ denotes Graph Retriever, we leverage the similarity between entities, relations, and the query to extract the initial subgraph. In the aligner module (Figure 1-➁), we use a GNN backbone to encode subgraph information. To tackle irrelevant knowledge noise and representation gaps, we introduce an LLM-based summarization technique with tailored prompts. By taking the query-answer pair and subgraph as input, the LLM generates a reasoning chain that links the query to the answer, identifying key intermediate nodes and bridging information gaps. Next, we propose the graph aligner to achieve dual alignment with this summarized reasoning chain: (1) Graph Node Alignment: A KL divergence loss [32] highlights critical nodes and reasoning edges, aligning node-level distributions and pruning irrelevant nodes. (2) Graph Representation Alignment: A contrastive loss with in-batch negative sampling aligns graph and language representations in a shared semantic space. Jointly optimizing these alignments enables effective knowledge pruning and unified representation learning. In the generation stage (Figure 1-➁), the aligned graph embedding is used as a graph token, concatenated with the query and pruned subgraph, and then fed into an LLM for answer generation. By leveraging the aligned graph, generator produces responses that reflect a deep understanding of graph’s structure and its relevance to query.
@@
Figure 1: The Align-GRAG framework includes three components: ➀ Graph Retriever, ➁ Graph Aligner, and ➂ Graph Generator. It first retrieve subgraphs, then trains the aligner with LLM-based summarized reasoning chain. The trained aligner prunes irrelevant nodes, generates graph embeddings, and feeds them into the LLM to produce the final answer.
3.1 Graph Retriever

### v2

@@ Abstract.
Large language models (LLMs) have demonstrated remarkable capabilities, but still struggle with issues like hallucinations and outdated information. Retrieval-augmented generation (RAG) addresses these issues by grounding LLM outputs in external knowledge with an Information Retrieval (IR) system. Building on this foundation, graph-based RAG systems go a step further by retrieving subgraphs, which preserve the relationships between knowledge entities and provide more comprehensive context. However, graph RAG faces two challenges: (1) Retrieving relevant information introduces irrelevant nodes (especially in dense graph databases, where retrieval usually extends to adjacent nodes), and leads to overly lengthy inputs that hinder efficiency; (2) The representation gap between graph and language during generation with LLMs limits the ability to fully leverage graph structures for enhanced understanding. To address these limitations, we propose Align-GRAG, a novel reasoning-guided dual alignment framework in post-retrieval phrase. It first formulates a subgraph by retrieving nodes and edges. Then an Aligner is proposed to jointly optimize a graph encoder with an LLM-summarized reasoning chain. It achieves dual alignment of graph node and representation by leveraging KL divergence loss and contrastive loss, facilitating efficient pruning of irrelevant knowledge and establishing a unified semantic space. The Generator integrates the aligned graph data with LLM to produce coherent and accurate answers. Experiments on the GraphQA benchmark across three tasks (including common sense reasoning, scene graph understanding, and knowledge graph reasoning) validate the effectiveness of our method. The codes are available in this repository [^1] https://anonymous.4open.science/r/Align-GRAG-F3D8/.
Retrieval-Augmented Generation, Large Language Model
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Computing methodologies Knowledge representation and reasoning
1. Introduction
@@
Recent advancements in natural language processing have demonstrated the remarkable capabilities of large language models (LLMs), such as GPT (OpenAI, 2024) and Llama (Touvron et al., 2023), in understanding, reasoning, and handling complex tasks (Zhao et al., 2023; Xu et al., 2024a). The massive parameterized knowledge encoded within LLMs has enabled them to excel in information retrieval (Zhu et al., 2023; AlKhamissi et al., 2022; Petroni et al., 2019). Despite these advancements, concerns about the interpretability of LLM remain, leading to challenges like hallucinations (fabrication of false or misleading information) (Huang et al., 2023; Xu et al., 2024c) and reliance on outdated data (Gao et al., 2023). These issues are especially concerning in high-stakes fields such as healthcare (He et al., 2023) and law (Lai et al., 2024).
@@
Retrieval-augmented generation (RAG) systems have been developed to solve these problems (Gao et al., 2023). By integrating external information retrieval (IR) systems, RAG retrieves knowledge from external databases, ensuring that LLMs are equipped with up-to-date and relevant information tailored to the user’s query. By grounding the LLM’s outputs in verifiable, external knowledge, RAG systems significantly improve the reliability and accuracy of content generated by LLMs (Fan et al., 2024). Nevertheless, in real-world scenarios, RAG systems often divide long context into independent chunks, overlooking the deeper connections between fragments and lacking a global perspective (Edge et al., 2024; Guo et al., 2024). Moreover, a significant portion of data inherently exhibits a graph-like structure, such as in recommendation systems (Guo et al., 2020), the Web, and knowledge graphs (Pan et al., 2024). To address these, researchers have proposed graph-based RAG (GRAG) (Edge et al., 2024; He et al., 2024; Tian et al., 2024; Sun et al., [n. d.]) as an enhancement to traditional RAG methods. Instead of retrieving isolated text chunks where information is treated independently of its structural relationships, GRAG retrieves subgraphs from databases (He et al., 2024). These subgraphs maintain the interconnected relationships between knowledge entities, providing richer context and enabling more accurate, contextually relevant responses.
Figure 1. Comparison of Align-GRAG with traditional RAG (including post-retrieval methods such as reranker). Left: query and textual graph (entities and relations). Middle (a): Traditional RAG retrieves top‑k passages and may use a reranker, but still suffers from noisy/irrelevant evidence and loses the underlying graph structure, which hurts answer grounding. Right (b): Align‑GRAG introduces a graph aligner that jointly performs node alignment and representation alignment, preserving structural relations and filtering irrelevant content, showcasing its effectiveness in graph RAG scenarios.
@@
However, despite its potential, integrating RAG with graphs is non-trivial, presenting two key challenges: (1) Irrelevant Knowledge: The retrieval stage often introduces irrelevant information, hindering the LLM’s attention on useful knowledge (especially in dense graph databases, where retrieval usually extends to adjacent nodes), and the generated inputs become overly lengthy, further causing computational inefficiency. As shown in Figure 1(a), some traditional RAG approaches (Xiao et al., 2024; Li et al., 2023b; Jia et al., 2024) aim to tackle this challenge by incorporating a reranker in the post-retrieval stage (Fan et al., 2024; Gao et al., 2023). For instance, BGE (Xiao et al., 2024) leverages cross-encoder-based architectures to effectively capture interactions between queries and documents, enabling more precise relevance estimation. However, in the GRAG setting, where chunks are typically represented as nodes or triplets, this approach can disrupt structural information and overlook the connections between knowledge. Meanwhile, existing GRAG methods (Guo et al., 2024) focus primarily on retrieval-stage optimizations (e.g., GraphRAG (Edge et al., 2024) builds community graphs for long documents to facilitate retrieval and summarization), resulting in relatively shallow similarity that lacks deeper semantic matching capabilities. (2) Representation Gap: In the generation stage, when integrating LLMs with graph structure embeddings, such as those generated by Graph Neural Network (GNN), a challenge arises due to the representation gap: graph embeddings are designed to capture structural properties, which misalign with the sequence-based representations used by LLMs. This misalignment limits the ability to fully leverage the graph for enhanced reasoning (Liu et al., 2023; Zhao et al., 2022). In the field of LLMs, most existing GRAG methods (Sun et al., [n. d.]; LUO et al., [n. d.]; Xu et al., 2025; Hu et al., 2022; Chen et al., 2024; Jiang et al., 2024) primarily treat graph data as plain text inputs, lacking a dedicated encoder for structural information, which limits their understanding capabilities. Some methods (Tian et al., 2024; Perozzi et al., 2024; He et al., 2024) incorporate graph encoders, but typically rely on straightforward approaches like concatenating graph embeddings with LLM inputs. For instance, G-Retriever (He et al., 2024) uses a projector to map graph embeddings, while GNP (Tian et al., 2024) applies cross-modality pooling to fuse text and graph representations. However, neither approach explicitly aligns these distinct representation spaces, leaving the underlying gap unresolved. These research gaps present a question: Can we better prune irrelevant nodes while keeping structural information, and effectively encode graph representations for utilization by LLMs?
@@
In this work, we propose Align-GRAG, a novel reasoning-guided dual alignment framework specifically designed for GRAG. During the retrieval stage, we calculate the similarity between nodes and edges to extract the initial subgraph. To address the two key challenges in GRAG, we introduce an innovative graph aligner module for the post-retrieval phase, as shown in Figure 1(b). We leverage LLM-based summarization ability with well-crafted prompts to generate reasoning chains that bridge information gaps. The aligner achieves dual alignment with reasoning chains as optimization target: (1) Graph Node Alignment, which ranks critical reasoning nodes, pruning irrelevant knowledge; and (2) Graph Representation Alignment, ensuring unified semantic space between graph and language representations via contrastive learning. This dual alignment refines subgraphs to focus on relevant knowledge, enabling the generator to produce accurate, context-aware responses grounded in the graph’s structure. Extensive experiments are conducted on GraphQA benchmark (He et al., 2024), covering tasks such as commonsense reasoning, scene graph understanding, and knowledge graph reasoning, which demonstrate the superior performance of Align-GRAG compared to other baselines and highlight the effectiveness of Aligner module. Overall, this work makes three key contributions:
@@
2. Related Work
2.1. Retrieval-Augmented Generation (RAG)
RAG (Fan et al., 2024; Gao et al., 2023; Fan et al., 2024) has been extensively studied to address the challenges of hallucination and outdated information (Huang et al., 2023; Xu et al., 2024c), which often lead to untrustworthy outputs from LLMs (Lewis et al., 2020). By integrating information retrieval systems, RAG has demonstrated its effectiveness in real-world applications, enhancing reliability of generated responses (Salemi and Zamani, 2024; Zamani and Bendersky, 2024; Yang et al., 2024; Sudhi et al., 2024). The RAG pipeline mainly includes: pre-retrieval (e.g., query rewriting and query expansion), retrieval, post-retrieval (e.g., reranking), and generation (Fan et al., 2024). Among these, reranking has emerged as a key technique for refining retrieval results. Initial retrieval (Robertson et al., 2009; Izacard et al., 2021) often relies on simple similarity scoring. Reranking (Li et al., 2023b; Xiao et al., 2024; Dong et al., 2024), on the other hand, employs more sophisticated models to reassess and reorder the initially retrieved documents based on their relevance to the query, thereby improving the quality and relevance of retrieved documents. For instance, bge-reranker (Xiao et al., 2024) leverages cross-encoders that perform full attention over the input pair, providing higher accuracy compared to embedding-based models. However, in GRAG scenarios, reranking may disrupt the inherent structural information of graph. Our proposed Align-GRAG effectively performs post-retrieval processing to extract relevant knowledge, while preserves the graph structural information.
2.2. Large Language Models on Graph
Graphs, composed of nodes and edges, are essential for modeling real-world relationships across various domains. In recent years, GNN (Shi et al., 2020; Veličković et al., 2018; Kipf and Welling, 2016; Wu et al., 2020) have emerged as a powerful tool for encoding graph structures. With the rise of LLMs and their demonstrated impressive capabilities, there is growing interest in integrating LLMs with graph learning techniques to enhance graph-based tasks (Huang et al., 2024b; Jin et al., 2024; Li et al., 2023a; Ren et al., 2024). Pioneering works (Tang et al., 2024; Perozzi et al., 2024; Chai et al., 2023) have incorporated GNNs by feeding graph tokens into LLMs. Beyond basic integration, recent studies (Qin et al., 2023; Zhu et al., 2024; Huang et al., 2024a) have focused on tighter architectural fusion, embedding graph neural layers within transformer architectures to enable seamless interaction between graph reasoning and natural language understanding . However, most existing researches (Ren et al., 2024) focus on tasks like node and graph classification. Our work addresses the challenges of retrieval-based graph QA, leveraging retrieval-augmented methods and graph learning techniques while overcoming the limitations of current frameworks.
2.3. Graph RAG
Traditional RAG often struggles to capture structured relational knowledge or global context. Graph Retrieval-Augmented Generation (GRAG) addresses these limitations by leveraging graph-structured knowledge for improved retrieval and reasoning (Baek et al., 2023; Wang et al., 2024; Han et al., 2024; Xu et al., 2024b; Zhang et al., 2025). GRAG utilizes graph databases (e.g., Freebase (Bollacker et al., 2008), Wikidata (Vrandečić and Krötzsch, 2014)) to retrieve graph elements like triples, paths, or subgraphs (Edge et al., 2024; Mavromatis and Karypis, 2024; Hu et al., 2024; Wu et al., 2024; Xiang et al., 2025; Xiao et al., 2025). Some approaches (Edge et al., 2024; Guo et al., 2024; Fan et al., 2025; Liang et al., 2024) focus on constructing large-scale graphs from text. For example, GraphRAG (Edge et al., 2024) uses LLMs to extract entities, partitions graphs using hierarchical clustering, and generates summaries to support retrieval. Other methods (Yasunaga et al., 2021; Jiang et al., 2022; He et al., 2024; Tian et al., 2024) encode graph data with GNNs and integrate it with language models for enhanced reasoning, while some translate natural language queries into logical forms for knowledge graph retrieval (Zhang et al., 2023; Yu et al., 2022; Luo et al., 2023; Xu et al., 2025; Hu et al., 2022). More recently, LLMs have been used as iterative agents for reasoning over knowledge graphs (Jiang et al., 2024; LUO et al., [n. d.]; Sun et al., [n. d.]; Chen et al., 2024; Zhang et al., 2024). However, existing GRAG approaches primarily emphasize optimizing retrieval but often lack effective post-retrieval strategies, leading to shallow subgraph-query similarity. To address this, our Align-GRAG introduces a novel dual alignment mechanism that bridges retriever and generator, improving overall performance.
3. Preliminaries
@@
This section presents relevant concepts and notations, including the definition of textual graphs and the task formalization.
@@
Textual Graph. A textual graph refers to a graph where the nodes and edges are supplemented with textual information, enabling the representation of both structural relationships and semantic content. Formally, a textual graph is denoted as $\mathcal{G}=(\mathcal{V},\mathcal{E},\{t_{n}\}_{n\in\mathcal{V}},\{t_{e}\}_{e\in\mathcal{E}})$, where, $\mathcal{V}$ and $\mathcal{E}$ represent the sets of nodes and edges, respectively. $t_{n}\in\mathcal{D}^{L_{n}}$ denotes the sequential text associated with a node $n\in\mathcal{V}$, where $\mathcal{D}$ is the vocabulary set, and $L_{n}$ is the length of the text associated with the node. Similarly, $t_{e}\in\mathcal{D}^{L_{e}}$ represents the sequential text associated with an edge $e\in\mathcal{E}$, and $L_{e}$ denotes the length of the text associated with the edge.
@@
Task Formulation. This work focuses on RAG with textual graph databases. The goal is to retrieve relevant information from the structured nature of textual graphs and generate coherent, contextually accurate responses. Formally, a given query $t_{q}=[x_{1},...,x_{n}]$, represented as a sequence of tokens from the vocabulary $\mathcal{D}$. The model retrieves relevant subgraphs $\mathcal{G}_{r}=(\mathcal{V}_{r},\mathcal{E}_{r})$ from $\mathcal{G}$ by computing the semantic similarity between the query $t_{q}$ and the node and edge texts $t_{n}$ and $t_{e}$. where $\mathcal{V}_{r}\subseteq\mathcal{V}$ and $\mathcal{E}_{r}\subseteq\mathcal{E}$. The retrieval process identifies relevant nodes, edges, and their associated texts in the context of the input question $t_{q}$. In the post-retrieval stage, we focus on refining the retrieved information to enhance the quality of input provided to LLM, ensuring more accurate and contextually relevant outputs. In the generation stage, the retrieved subgraph, along with the query $t_{q}$ and prompt $\mathcal{P}$, is provided as input to LLM. The LLM generates the final output $Y=[y_{1},...,y_{m}]$, grounded in the retrieved knowledge and the contextual understanding of the graph structure.
4. Methodology
@@
In this section, we introduce the Align-GRAG framework, which is composed of three components: Graph Retriever (Sec. 4.1), Graph Aligner (Sec. 4.2), and Graph Generator (Sec. 4.3) modules. The overall framework pipeline is depicted in Figure 2. In Figure 1, ➀ denotes Graph Retriever, we leverage the similarity between entities, relations, and the query to extract the initial subgraph. In the aligner module (Figure 2-➁), we use a GNN backbone to encode subgraph information. To tackle irrelevant knowledge noise and representation gaps, we introduce an LLM-based summarization technique with tailored prompts. By taking the query-answer pair and subgraph as input, the LLM generates a reasoning chain that links the query to the answer, identifying key intermediate nodes and bridging information gaps. Next, we propose the graph aligner to achieve dual alignment with this summarized reasoning chain: (1) Graph Node Alignment: A KL divergence loss (Kullback and Leibler, 1951) is utilized to highlight critical nodes and reasoning edges, aligning node-level distributions and pruning irrelevant nodes. (2) Graph Representation Alignment: A contrastive loss with in-batch negative sampling is further designed to align graph and language representations in a shared semantic space. Jointly optimizing these alignments enables effective knowledge pruning and unified representation learning. In the generation stage (Figure 2-➁), the aligned graph embedding is used as a graph token, concatenated with the query and pruned subgraph, and then fed into an LLM for answer generation. By leveraging the aligned graph, generator produces responses that reflect a deep understanding of graph’s structure and its relevance to query.
@@
Figure 2. The Align-GRAG framework includes three components: ➀ Graph Retriever, ➁ Graph Aligner, and ➂ Graph Generator. It first retrieves subgraphs, then trains the aligner with an LLM-based summarized reasoning chain. The trained aligner prunes irrelevant nodes, generates graph embeddings, and feeds them into the LLM to produce the final answer.
4.1. Graph Retriever